OpenStudy (anonymous):

what is the slope of the line that passes through the points (-2,4) and (8,-1)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

start by finding the slope of the line then using the slope of the line and one of the points, plug them into point-slope form

OpenStudy (blurbendy):

do you understand @Sarahneedshelp ?

OpenStudy (anonymous):

@blurbendy no and i'm about to fail the regents and i have so many questions on this homework left #pleasehelp

OpenStudy (blurbendy):

what he is basically saying is that we need to find the slope first. slope = y2 - y1 / x2 - x1 so let's substitute those values in with the points we have (-1 -4) / (8 + 2) = -5/10 = -1/2 do you see how i got that?

OpenStudy (anonymous):

@blurbendy yes actually omg, thank you can you help me with the rest of my homework?

OpenStudy (blurbendy):

sure, no problem

OpenStudy (anonymous):

@blurbendy you're actually a lifesaver the next question i don't get is what is the equation of the line that is perpendicular to the line y-2x=4 and passes through point (2,4)

OpenStudy (blurbendy):

okay, first we need to get the equation in the form y = mx + b, so y - 2x = 4 y = 2x + 4 now we can see that the slope is 2 however we need to the find the equation of the line that is perpendicular, and to do that we take the OPPOSITE reciprocal of the slope (2), which is -1/2 follow so far?

OpenStudy (anonymous):

so the slope would be -1/2 and the equation would be y=-1/2x+5

OpenStudy (blurbendy):

perfect, you got 5 by substituting (2, 4) in right?

OpenStudy (anonymous):

@blurbendy yes! i just plugged them in

OpenStudy (blurbendy):

good job

OpenStudy (anonymous):

@blurbendy thank ya:) okay next question the diagonal of a square is 8. What is the area of a square?

OpenStudy (blurbendy):

Area of a square using its diagonal = 1/2 x d^2 can you solve from here?

OpenStudy (blurbendy):

the 'x' is times, not a variable

OpenStudy (anonymous):

is the answer 32? or am i way off?

OpenStudy (blurbendy):

that's absolutely right!

OpenStudy (anonymous):

oh my god yes well now we got a proof....

OpenStudy (blurbendy):

:o

OpenStudy (anonymous):

Given: rhombus RSTV, VTX, STW, SX, VW and <RSX=<RVW Prove TX=TW

OpenStudy (blurbendy):

hmm, do you mind bumping your question? someone with better proof skills will be able to help you more than I can.

OpenStudy (anonymous):

How do i do that?:)

OpenStudy (blurbendy):

go to the top where you had your first question. it should give you the option to bump

OpenStudy (blurbendy):

if you don't see it, then just make a new question for it

OpenStudy (blurbendy):

good luck!

OpenStudy (anonymous):

@blurbendy can you help me with the rest of my hw?

OpenStudy (blurbendy):

ill try!

OpenStudy (blurbendy):

so, no more?